OpenStudy (anonymous):

If f(x) = -2x + 1, find f(-6

OpenStudy (haseeb96):

just put -6 in the place of x in the f(x)=-2x +1 and solve it then get the answer

OpenStudy (haseeb96):

@kaylamarie081 put -6 and show me yur work

OpenStudy (anonymous):

Okay, give me a minute to solve

OpenStudy (haseeb96):

f(-6)=-2(-6)+1 = 12 +1 = 13 is the correct answer

OpenStudy (anonymous):

Or nottt. :o thank you.

OpenStudy (haseeb96):

if u think that i have help u then click on best response

OpenStudy (princeharryyy):

oh lol!!! people help for getting medals only hehe :D